### 01 Feb 2024 v1

@@ Appendix A Msss matyices
The CP-even neutral scalars have the composition $S^{T}=(h_{d},h_{u},(\tilde{\nu}_{i}^{c})^{\Re})$, one can write the mass matrix $M_{S}^{2}$ as:
$M_{S}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}h_{d}}^{2}&M_{h_{d}h_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{h_{u}h_{d}}^{2}&M_{h_{u}h_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\Re}h_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}h_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{1}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{2}^{c})^{\Re}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\Re}(\tilde{\nu}_{3}^{c})^{\Re}}^{2}\\ \end{array}\right)$ (46)
$M_{h_{d}h_{d}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{d}^{2}+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (47)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{11},$
$M_{h_{d}h_{u}}^{2}\!\!\!\!\!$ $=-{{G^{2}}\over{2}}v_{d}v_{u}-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}+2|\lambda_{i}|^{2}v_{d}v_{u}$ (48)
$-\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{12},$
$M_{h_{u}h_{u}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{u}^{2}+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (49)
$+\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{22},$
$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+2\lambda_{i}\lambda_{j}^{*}v_{d}\upsilon_{\nu_{j}^{c}}$ (50)
$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{j}^{c}}+\Delta_{1(2+i)},$
$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=-\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+2\lambda_{i}\lambda_{j}^{*}v_{u}\upsilon_{\nu_{j}^{c}}$ (51)
$-2\Re(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{j}^{c}}+\Delta_{2(2+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{i}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=(A_{\kappa}\kappa)_{ijk}\Re(e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}})\upsilon_{\nu_{k}^{c}}+\Delta_{(2+i)(2+j)}$ (52)
$+\Re(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{j}^{c}}}},$
$M_{(\tilde{\nu}_{i}^{c})^{\Re}(\tilde{\nu}_{j}^{c})^{\Re}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{d}^{2}+v_{u}^{2})+\Delta_{(2+i)(2+j)}.$ (53)
@@
The rediative corrections from the top quark and bottom quark and their corresponding supersymmetric partners, the corrections in the mass matrix can be expressed as:
@@
$\Delta_{1(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (55)
×$\times{{|\mu|^{2}\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{2(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\Bigg{\}}$ (56)
×$\times\Bigg{\{}\ln{{m_{\tilde{t}_{1}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{|A_{t}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{(2+i)(2+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}$ (57)
@@
$\Delta_{11}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}m_{\tilde{b}_{2}}^{2}}\over{m_{b}^{4}}}+{{2|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}$ (58)
$+{{(|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{12}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (59)
@@
×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{22}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{(|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),\Bigg{\}}$ (60)
@@
$\Delta_{1(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (61)
@@
×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$

### 06 Mar 2024 v2

@@ Acknowledgements.
The $CP$-even neutral scalars have the composition $S^{T}=(h_{d},h_{u},(\tilde{\nu}_{i}^{c})^{\text{Re}})$, and one can write the mass matrix $M_{S}^{2}$ as
$M_{S}^{2}=\left(\begin{array}[]{ccccc}M_{h_{d}h_{d}}^{2}&M_{h_{d}h_{u}}^{2}&M_{h_{d}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{h_{d}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{h_{d}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{h_{u}h_{d}}^{2}&M_{h_{u}h_{u}}^{2}&M_{h_{u}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{h_{u}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{h_{u}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}h_{d}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}h_{u}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{1}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}h_{d}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}h_{u}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{2}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}h_{d}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}h_{u}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{1}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{2}^{c})^{\text{Re}}}^{2}&M_{(\tilde{\nu}_{3}^{c})^{\text{Re}}(\tilde{\nu}_{3}^{c})^{\text{Re}}}^{2}\\ \end{array}\right)$ (46)
$M_{h_{d}h_{d}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{d}^{2}+\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\tan\beta$ (47)
$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\tan\beta+\Delta_{11},$
$M_{h_{d}h_{u}}^{2}\!\!\!\!\!$ $=-{{G^{2}}\over{2}}v_{d}v_{u}-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}+2|\lambda_{i}|^{2}v_{d}v_{u}$ (48)
$-\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}+\Delta_{12},$
$M_{h_{u}h_{u}}^{2}\!\!\!\!\!$ $={{G^{2}}\over{2}}v_{u}^{2}+\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}\upsilon_{\nu_{i}^{c}}\cot\beta$ (49)
$+\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})\upsilon_{\nu_{i}^{c}}\upsilon_{\nu_{j}^{c}}\cot\beta+\Delta_{22},$
$M_{h_{d}(\tilde{\nu}_{i}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{u}+2\lambda_{i}\lambda_{j}^{*}v_{d}\upsilon_{\nu_{j}^{c}}$ (50)
$-2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{u}\upsilon_{\nu_{j}^{c}}+\Delta_{1(2+i)},$
$M_{h_{u}(\tilde{\nu}_{i}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=-\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}v_{d}+2\lambda_{i}\lambda_{j}^{*}v_{u}\upsilon_{\nu_{j}^{c}}$ (51)
$-2\text{Re}(e^{-i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}}e^{i\phi_{\upsilon_{\nu_{j}^{c}}}}\lambda_{k}^{\ast}\kappa_{ijk})v_{d}\upsilon_{\nu_{j}^{c}}+\Delta_{2(2+i)},$
$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{i}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=(A_{\kappa}\kappa)_{ijk}\text{Re}(e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}}e^{i\phi_{\upsilon_{\nu_{k}^{c}}}})\upsilon_{\nu_{k}^{c}}+\Delta_{(2+i)(2+j)}$ (52)
$+\text{Re}(e^{i\phi_{\lambda}}e^{-i\phi_{v_{u}}}e^{i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})(A_{\lambda}\lambda)_{i}{{v_{d}v_{u}}\over{\upsilon_{\nu_{j}^{c}}}},$
$M_{(\tilde{\nu}_{i}^{c})^{\text{Re}}(\tilde{\nu}_{j}^{c})^{\text{Re}}}^{2}\!\!\!\!\!$ $=|\lambda_{i}|^{2}(v_{d}^{2}+v_{u}^{2})+\Delta_{(2+i)(2+j)}.$ (53)
@@
The radiative corrections from the top quark and bottom quark and their corresponding supersymmetric partners, the corrections in the mass matrix, can be expressed as
@@
$\Delta_{1(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (55)
×$\times{{|\mu|^{2}\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{2(2+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}\Bigg{\}}$ (56)
×$\times\Bigg{\{}\ln{{m_{\tilde{t}_{1}}^{2}}\over{m_{\tilde{t}_{2}}^{2}}}+{{|A_{t}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})\Bigg{\}},$
$\Delta_{(2+i)(2+j)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}\Bigg{\{}{{({1\over 2}\upsilon_{\nu_{j}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\cot\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d})^{2}}\over{(m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2})^{2}}}$ (57)
@@
$\Delta_{11}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}m_{\tilde{b}_{2}}^{2}}\over{m_{b}^{4}}}+{{2|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}$ (58)
$+{{(|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{12}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (59)
@@
×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{22}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{(|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu)^{2}}\over{(m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2})^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),\Bigg{\}}$ (60)
@@
$\Delta_{1(2+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}\Bigg{\{}{{{1\over 2}\upsilon_{\nu_{j}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$ (61)
@@
×$\times\Bigg{\{}\ln{{m_{\tilde{b}_{1}}^{2}}\over{m_{\tilde{b}_{2}}^{2}}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$